OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5. 6 -6 All real numbers No solution

HanAkoSolo (jamierox4ev3r):

just distribute the value in this problem: -3(x+3) = -3x-9

OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4. 1 -1 -5 -25

OpenStudy (anonymous):

3(x + 1) = -2(x - 1) - 4 3(x + 1) = -2x + 2 - 4 3x + 3 = -2x + 2 - 4 3x-2x=2-4-3 x=-5

OpenStudy (anonymous):

Understand?